Circular No. 7463

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000cx IN NGC 524 R. Chornock, D. C. Leonard, A. V. Filippenko, and W. D. Li, Uni-

versity of California, Berkeley; and E. L. Gates and K. Chloros, Lick Observatory, write: “Inspection of a CCD spectrum (range 390–720 nm) of SN 2000cx (cf. IAUC 7458) obtained on July 23 UT with the Nickel 1-m reflector at Lick Observatory reveals that the object is a peculiar type-Ia supernova, strongly resembling the overluminous SN 1991T a few days before maximum brightness (e.g., Filippenko et al. 1991, Ap.J. 384, L15). The Fe III absorption troughs near 430 and 490 nm are prominent, but Si II at 612 nm is weak. The object is in the outskirts of the host galaxy, which has a Hubble type of S0; this is the first time that a SN 1991T-like type-Ia object has been seen in such an early-type galaxy, in a region devoid of obvious dust and massive stars. Frequent observations at all wavelengths are encouraged.”

K. Krisciunas, University of Washington (UW), reports preliminary photometry of SN 2000cx, reduced by him from images taken by C. Logan and T. Tavenner (UW) using the Manastash Ridge Observatory 0.76-m reflector near Ellensburg, WA: July 19.42 UT, V = 14.00, B − V = +0.13, V − R = +0.01, V − I = −0.03 (uncertainties ± 0.04 mag); July 20.43, V = 13.82, B − V = +0.16, V − R = 0.00, V − I = −0.07 (uncertainties ± 0.02 mag).

XTE J1859+226 P. Rodriguez-Gil and J. Casares, Instituto de Astrofisica de Canarias

(IAC), report on optical photometry of this soft x-ray transient (IAUC 7274, 7279): “R-band photometry of this x-ray transient has been recently performed with the 1.0-m Jacobus Kapteyn Telescope on La Palma and the 0.8-m IAC-80 telescope on Tenerife. After the minioutburst reported on IAUC 7451, the brightness is steeply decreasing at a rate of ∼ 0.4 mag/day. Measured magnitudes: July 17.10 UT, R = 19.1; 19.92, 19.9; 21.94, 20.9; 22.93, 21.6. Further coverage is urged both now (while in decline) and during the following months, since a new minioutburst may occur.”

SUPERNOVA 2000cw IN MCG +5-56-007 Corrigendum. On IAUC 7457, lines 1–2, for on behalf a collabora-

tion read on behalf of a collaboration known as the ‘Network of European Observatories in the North (NEON) Observation Euroschool’,

2000 July 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green